Suppose the risk-free rate is 2.75% and an analyst assumes a market risk premium of 5.07%. Firm A just paid a dividend of $1.2 decimal places
Answer #1

Value of Firm A=Number of shares*D0*(1+g)/(rf+b*MRP-g)=268*1.24*1.0473/(2.75%+1.43*5.07%-4.73%)=6604.025275 million